OpenStudy (anonymous):

the diagonal of a square has a length of 10cm. find the length of a side of a square.

OpenStudy (anonymous):

let the length of the side be 'a'.  consider the triangle ADC by applying Pythagoras theorem you get, 10^2=a^2+a^2 =>100=2a^2 =>a^2=50 =>a=sqrt(50) =>a=5 x root (2)